OpenStudy (anonymous):

Find all zeros and their multiplicities. f(x) = 7x^3+x

OpenStudy (campbell_st):

so first let f(x) = 0 so then \[7x^3 + x = 0\] factoring gives \[x(7x^2 + 1) = 0\] so 1 zero occurs when x = 0 to find the other zeros you need to solve \[7x^2 + 1 = 0\] you may need the general quadratic formula...

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

\[-b \pm \sqrt{b ^{2}-4ac}/2a\]

OpenStudy (anonymous):

how would i do that?

OpenStudy (campbell_st):

well b = 0, a = 7 and c = 1 so \[x = \frac{0 \pm \sqrt{0^2 - 4 \times 7 \times 1}}{2 \times 7}\] which will give a complex root... and each root has a multiplicity of 1

OpenStudy (anonymous):

okie dokes got it. thank you again!